The sample includes 55% of women and 45% of men. It is known that among them 35% of women and 65% of men smoke. Person X is known to be a non-smoker. Find the probability that X is male. Round your answer to the nearest hundredth
1
Expert's answer
2021-10-06T16:29:20-0400

A - "Person is a non-smoker"

B - "Person is male"

"P(B|A) = {\frac {P(A⋂B)} {P(A)}} = {\frac {0.45*0.35} {0.55*0.65+0.45*0.35}} = {\frac {0.1575} {0.515}} = 0.31"